Simplify the expressions and find the value if x is equal to 2 
x + 7 + 4 (x – 5)

Grade:12

4 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that x = 2
We have,
= x + 7 + 4x–20
= 5x + 7–20
Then, substitute the value of x in the equation
= (5 × 2) + 7–20
= 10 + 7–20
= 17–20
=-3

Thanks
Jitendra Kumar
3 years ago
Dear Student
 
     x + 7 + 4( x-5)
 =  x + 7 + 4x - 4 × 5
 =  x + 7 + 4x - 20
 =  x + 4x +7 - 20
 = 5x - 13
Now, we find the valu of the expression at x = 2
So we put x=2 
  = (5 × 2) - 13
  = 10 - 13
  = -3
 
Thanks
 
 
Ram Kushwah
3 years ago
Dear The given expression is:
x+7+4(x-5)
 
=x+7+4x-20
=5x-13
Given that x=2
The value=5*2-13
=10-13
 
=-3
Shraddha Singh
2 years ago
Here, the given expression is 
x+7+4(x-5)
 
At x = 2,
the expression becomes,
2+7+4(2-5)
=9+4(-3)
= 9-12
= -3
 
Hope it helps!!
Thank you